# MIRROR: Multi-Modal Pathological Self-Supervised Representation Learning via Modality Alignment and Retention

### v3

@@ Abstract
Index Terms:
Pathology, Whole Slide Image (WSI), Transcriptomics, Self-Supervised Learning (SSL), Multimodal Learning
@@ IV-E Qualitative Analysis
TABLE VIII: Ablation study of loss term weights.
Module Setting Performance
$\lambda_{\alpha}$ $\lambda_{\beta}$ $\lambda_{\gamma}$ Subtyping Survival
0.3 0.5 0.2 0.942 ($\pm$ 0.020) 0.650 ($\pm$ 0.053)
0.3 0.2 0.5 0.950 ($\pm$ 0.021) 0.657 ($\pm$ 0.041)
0.5 0.3 0.2 0.955 ($\pm$ 0.023) 0.665 ($\pm$ 0.082)
@@ V-D Computational Complexity
We further analyzed the computational cost of MIRROR to assess the trade-offs introduced by the proposed architecture and modules. The number of parameters and floating-point operations (FLOPs) per sample were measured and reported.
When applied to downstream tasks, where a single linear classifier is attached to the backbone, MIRROR comprises 33.1033.10M parameters and requires 90.4490.44G FLOPs per forward pass. During pretraining, the complexity increases to 48.9748.97M parameters and 139.14139.14G FLOPs, reflecting the additional overhead of these components. In comparison, TANGLE remains considerably lighter in both pretraining and downstream evaluation, with 11.4111.41M parameters and 9.699.69G FLOPs.
@@ V-F Limitations and Future Work
Method TCGA-NSCLC TCGA-RCC TCGA-COADREAD
Acc F1 Acc F1 Acc F1
TANGLE 0.947 0.987 0.956 0.908 0.744 0.448
(± $\pm\,0.044$) (± $\pm\,0.110$) (± $\pm\,0.004$) (± $\pm\,0.148$) (± $\pm\,0.175$) (± $\pm\,0.020$)
MIRROR 0.993 0.993 0.983 0.914 0.822 0.523
(± $\pm\,0.006$) (± $\pm\,0.006$) (± $\pm\,0.012$) (± $\pm\,0.145$) (± $\pm\,0.050$) (± $\pm\,0.075$)
@@
Method TCGA-NSCLC TCGA-RCC TCGA-COADREAD
TANGLE 0.536 (± $\pm\,0.025$) 0.744 (± $\pm\,0.130$) 0.511 (± $\pm\,0.016$)
MIRROR 0.602 (± $\pm\,0.044$) 0.778 (± $\pm\,0.055$) 0.693 (± $\pm\,0.045$)

### v4

@@ V-B Adaptability Across Disease Types
TABLE VIII: Ablation study of loss term weights.
Module Setting Performance
$\lambda_{\alpha}$ $\lambda_{\beta}$ $\lambda_{\gamma}$ Subtyping Survival
0.3 0.5 0.2 0.942 ($\pm$ 0.020) 0.650 ($\pm$ 0.053)
0.3 0.2 0.5 0.950 ($\pm$ 0.021) 0.657 ($\pm$ 0.041)
0.5 0.3 0.2 0.955 ($\pm$ 0.023) 0.665 ($\pm$ 0.082)
@@
To further evaluate generalizability beyond TCGA cohorts, we conduct an external experiment on the Clinical Proteomic Tumor Analysis Consortium Kidney cohort (CPTAC-KIDNEY) [86], which contains 783 slides of clear cell renal cell carcinoma (ccRCC) and 126 slides of non–clear cell renal cell carcinoma (non-ccRCC). The evaluation follows a 5-fold cross-validation protocol consistent with our downstream settings. In contrast to the TCGA-RCC dataset, which is larger and relatively more balanced across subtypes with 519 slides for KIRC, 300 slides for KIRP, and 121 slides for KICH, the CPTAC-KIDNEY is smaller in scale and exhibits a markedly imbalanced subtype distribution, with ccRCC substantially outnumbering non-ccRCC. Despite these challenges, MIRROR maintains strong transfer performance, outperforming TANGLE on both subtype classification and survival analysis tasks. These results demonstrate MIRROR’s robustness to distribution shift across smaller and more heterogeneous datasets.
@@ V-C1 Sensitivity to Input Perturbations
V-C2 Sensitivity to Distribution Shifts in Survival Data
In addition to the local sensitivity analysis, we further examine each model’s sensitivity to distributional factors in the survival analysis task, namely label imbalance and right-censoring, using the TCGA-NSCLC cohort. Following standard practice, survival times are discretized into four bins with sample counts of $190$, $202$, $280$, and $250$, totaling $922$ cases, among which $563$ ($61.1\%$) are censored. To evaluate robustness under these conditions, we compute three censoring-aware metrics on the training folds. The Integrated Brier Score (IBS) reflects the time-integrated prediction error, the mean time-dependent Area Under the Receiver Operating Characteristic Curve (t-AUC) summarizes discriminative ability across time bins, and the mean Brier score represents the average bin-wise squared error. Lower IBS and Brier values together with higher t-AUC values indicate better performance. MIRROR consistently outperforms TANGLE across all three measures, achieving an IBS of $0.119$ versus $0.217$, a mean t-AUC of $0.940$ versus $0.598$, and a mean Brier score of $0.141$ versus $0.230$. These results demonstrate that MIRROR attains stronger discriminative power and lower prediction error despite substantial censoring and label imbalance in the cohort.
@@ V-D Computational Complexity
We further analyzed the computational cost of MIRROR to assess the trade-offs introduced by the proposed architecture and modules. The number of parameters, floating-point operations (FLOPs) per sample, runtime throughput, and memory usage were measured and reported to provide a comprehensive evaluation.
When applied to downstream tasks, where a single linear classifier is attached to the backbone, MIRROR comprises 33.1033.10M parameters and requires 90.4490.44G FLOPs per forward pass. During pretraining, the complexity increases to 48.9748.97M parameters and 139.14139.14G FLOPs, reflecting the additional overhead of these components. In comparison, TANGLE remains considerably lighter in both pretraining and downstream evaluation, with 11.4111.41M parameters and 9.699.69G FLOPs. For runtime efficiency, we further measured throughput and peak GPU memory usage on our experimental environment. At inference with a batch size of 16, MIRROR processes $227.88$ samples/s, whereas TANGLE reaches 2,593.772{,}593.77 samples/s. During pretraining, also with a batch size of 16, MIRROR attains $61.58$ samples/s per optimization step, while TANGLE achieves $799.19$ samples/s. The peak GPU memory consumption during training is $10.30$ GB for the MIRROR , $19.24$ GB for the pretraining model, and $3.17$ GB for TANGLE.
@@ V-E Robustness to Domain Shift
Method TCGA- NSCLC TCGA- RCC TCGA- COADREAD CPTAC- KIDNEY
Acc F1 Acc F1 Acc F1 Acc F1
TANGLE 0.947 0.987 0.956 0.908 0.744 0.448 0.934 0.827
(± $\pm\,0.044$) (± $\pm\,0.110$) (± $\pm\,0.004$) (± $\pm\,0.148$) (± $\pm\,0.175$) (± $\pm\,0.020$) (± $\pm\,0.005$) (± $\pm\,0.011$)
MIRROR 0.993 0.993 0.983 0.914 0.822 0.523 0.986 0.964
(± $\pm\,0.006$) (± $\pm\,0.006$) (± $\pm\,0.012$) (± $\pm\,0.145$) (± $\pm\,0.050$) (± $\pm\,0.075$) (± $\pm\,0.001$) (± $\pm\,0.002$)
@@ V-F Limitations and Future Work
Method TCGA- NSCLC TCGA- RCC TCGA- COADREAD CPTAC- KIDNEY
TANGLE 0.536 (± $\pm\,0.025$) 0.744 (± $\pm\,0.130$) 0.511 (± $\pm\,0.016$) 0.614 (± $\pm\,0.087$)
MIRROR 0.602 (± $\pm\,0.044$) 0.778 (± $\pm\,0.055$) 0.693 (± $\pm\,0.045$) 0.816 (± $\pm\,0.060$)